write an equation of the line through(3,-2_ and parallel to 2x-y+5=0.

Answer by nerdybill(7384)

write an equation of the line through(3,-2) and parallel to 2x-y+5=0.
.
find slope of given line by putting into "slope-intercept" form:
2x-y+5=0
2x+5=y
slope is 2
a line that is parallel will have the same slope.
.
Our new line has slope of 2 and crossing (3,-2). Put into "point-slope" form:
y - y1 = m(x - x1)
y - (-2) = 2(x - 3)
y + 2 = 2x - 6
y = 2x - 8 (answer in "slope-intercept" form)